+ 12 points Ser.P10 10 PO MI F My Note An automobile tire is inflated with air originally at 10.0°C and normal atmospheric pressure. During the process, the air is compressed to 24.0% of its original volume and the temperature is increased to 47,0C (a) What is the tire pressure in pa scals? Pa (b) After the car is driven at high speed, the tires air temperature rises to 85.0°C and the tires interior volume increases by J.00%. What is the new tire pressure (absolute) in pascals? Pa Need Help? E Read. Master.
Answer #1

a)

We can use the combined gas law:

P1*V1/T1 = P2*V2/T2

We know that: V2 = 0.24 V1

P1 = 1 atm = 101325 Pa

T1 = 10 C = 10+273 = 283 K
T2 = 47 C = 47 + 273 = 320 K

(101325)*V1/(283) = P2*(0.24V1)/320

We then get:
P2 = (320)*(101325)/(283*0.24)

= 477385.159011 Pascals

(b)

Again we have the combined gas law with:

It's because it increased by 3% so,

V2 = 1.03V1

P1 =477385.159011 Pascals

T1 = 47 C = 320 K
T2 = 85 C = 358 K

(477385.159011)(V1)/(320) = (P2)*(1.03V1)/(358)

P2 = (477385.159011)*(358)/(320)*(1.03)

P2 = 518519.074411 Pascals

I hope help you !!